OpenStudy (anonymous):

find the derivative of f(x)=x^2 ln x

OpenStudy (anonymous):

\[(fg)'-f'g+g'f\] with \[f(x)=x^2,f'(x)=2x,g(x)=\ln(x), g'(x)=\frac{1}{x}\]

OpenStudy (anonymous):

.y = (lnx)/x^2 ------------------------------- y = f(x)/g(x) dy/dx = (f' *g - g' *f)/g^2 -------------------------- dy/dx = [(1/x) *x^2 - (2x) lnx] / x^4 dy/dx = [x - (2x) lnx] / x^4 dy/dx = [1 - 2 lnx] / x^3 >>>> ---------- d^y/dx^2 = [-(2/x)*x^3 - 3x^2(1 - 2 lnx)] / x^6 d^y/dx^2 = - [2 + 3(1 - 2 lnx)] / x^4 d^y/dx^2 = - [5 -6 lnx)] / x^4 d^y/dx^2 = [- 5 +6 lnx] / x^4 >>>

OpenStudy (anonymous):

so f'(x)=x+2x lnx

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

wy